News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 20 July 2007 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Chandra passed through the final eclipse of the 2007 Summer season on July 12 with nominal power and thermal performance. A real-time procedure was conducted on July 13 set the short eclipse flag to 'false'. A second real-time procedure was executed on July 13 to clear a glitch counter that incremented during eclipse exit. The counter increment is expected under certain timing and thermal conditions.

A commanding anomaly was observed on July 15 when the the first block of commands sent during a pass to turn on the downlink signal was resent autonomously by DSN station DSS-46. There were no impacts from the duplicate commanding however we are working closely with DSN to understand the cause of the anomaly.

The targets recommended for acceptance from the Cycle-9 Peer review were posted on schedule on July 13. For details see: http://cxc.harvard.edu/target_lists/

The schedule of targets for the next week is shown below and includes an observation of PSR B1259-63 coordinated with VLBI observations.

------------------------------------------
Radiation Belts                     Jul 23
NGC 6826               ACIS-S
Vega (4 obs)           HRC-S        Jul 24
Vega                   HRC-I
GC-1                   HRC-I
GC-2                   HRC-I
GC-4                   HRC-I
GC-5                   HRC-I
GC-6                   HRC-I
GC-7                   HRC-I
RXJ1131-1231           ACIS-S
SDSS J1212+5341        ACIS-S
RXJ 1257+4738          ACIS-S
AX J185750+0240        ACIS-S       Jul 25
Abell 2395             ACIS-S
IC 5325                ACIS-S
Radiation Belts
3EG J1826-1302         ACIS-I       Jul 26
J1826-14               ACIS-I       Jul 27
Abell 2244             ACIS-I
Abell 222              ACIS-I
Orion Trapezium Cl     ACIS-S/HETG
PSR B1259-63           ACIS-S/HETG  Jul 28
Orion Trapezium Cl     ACIS-S/HETG
Radiation Belts
IRAC Dark Field        ACIS-I       Jul 29
A1446                  ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports